OpenStudy (anonymous):

A gem store sells beads made of amber and quartz. For 4 amber beads and 4 quartz beads, the cost is $46.00. For 1 amber bead and 3 quartz beads, the cost is $14.50. Find the price of each type of bead.

OpenStudy (anonymous):

4a+4q=46 a+3q=14.5 Solve the system. Where a - amber beads; q - quartz beads.

OpenStudy (anonymous):

i need more help; okay so do i solve each seperatly?

OpenStudy (anonymous):

This is a system of equations, it can't be solved separately. You need to use techniques you have learned at school to solve them. One of the techniques is to try to eliminate say a. To do that you can multiply the second equation by (-4). 4a+4q=46 -4a-12q=-58 Now you add both systems and you cancel the a because 4a-4a=0. And you get -8q=-12 Can you continue from here and solve the whole system?

OpenStudy (anonymous):

Have you solved this yet?

OpenStudy (anonymous):

Ok your soln is q=$1.5, a=$10. Have fun lol